OpenStudy (anonymous):

Out of 36 students in a hostel, 19 left. Due to this, the expenditure of the mess decreased by $48 per day, while the average expenditure per head increased by $5. What is the sum of original expenditure and the new expenditure per day of the mess?

OpenStudy (anonymous):

$456 $456 $456 $456 $456 (B)

OpenStudy (anonymous):

metal

OpenStudy (dan815):

48/19 = avg expenditure before 19 ppl left and add 5 to that multiply by the number of ppl left

OpenStudy (anonymous):

2.52+5*19=97.52

OpenStudy (anonymous):

36(x+5)=17x-48, 36x+180=17x-48 x=6.9

OpenStudy (anonymous):

36*6.9=252.10 is the original exp 17*(5+6.9)=202.3is the new exp so 452.3 is the ans i think